# Generalization Bounds via Meta-Learned Model Representations: PAC-Bayes and Sample Compression Hypernetworks

### v2

@@ 1 Introduction
Machine learning is increasingly shaping our daily lives, a trend accelerated by rapid advancements in deep learning and the widespread deployment of these models across various applications (e.g., language models and AI agents). Ensuring the reliability of machine learning models is therefore more critical than ever. A fundamental aspect of reliability is understanding how well a model generalizes beyond its training data, particularly for modern neural networks, whose complexity makes it challenging to obtain strong theoretical guarantees. One way to assess generalization is through probabilistic bounds on a model’s error rate. Applying these techniques to deep neural networks is challenging because classical approaches struggle to account for the true effective complexity of such models (Zhang et al., 2017), which may not be well captured by naive measures such as parameter count (Belkin et al., 2019). However, prior work suggests that if one can obtain a more compact yet expressive representation of a model’s complexity, tighter generalization bounds are possible (Dziugaite et al., 2020; Wang et al., 2022; Kawaguchi et al., 2023; Lotfi et al., 2022).
@@
Related works. The meta-learning framework was pioneered by Baxter (2000), where a learning problem encompasses multiple tasks under the assumption that all learning tasks are independently and identically distributed (i.i.d.) from a task environment distribution. Various standpoints have been considered to derive generalization guarantees in this setting, for example the VC-theory (Maurer, 2009; Maurer et al., 2016) and algorithmic stability (Maurer, 2005; Chen et al., 2020). In recent years, the PAC-Bayesian framework has been the foundation of a rich line of work: Pentina & Lampert (2014); Amit & Meir (2018); Rothfuss et al. (2021); Liu et al. (2021); Guan et al. (2022); Guan & Lu (2022); Rezazadeh (2022); Zakerinia et al. (2024). Up to our knowledge, using the sample compression framework (Littlestone & Warmuth, 1986) to derive generalization bounds in meta-learning is an idea that has not been explored yet.
@@ 2 Learning Theory and Generalization Bounds
The prediction problem. A dataset $S=\{({\mathbf{x}}_{j},y_{j})\}_{j=1}^{m}$ is a collection of $m$ examples, each of them being a feature-target pair $(\mathbf{x},y)\in\mathcal{X}\times\mathcal{Y}$, and a predictor is a function $h:\mathcal{X}\rightarrow\mathcal{Y}$. We denote $\operatorname{\mathcal{H}}$ the predictor space. Let $A$ be a learning algorithm $A:\bigcup_{k\in\mathbb{N}}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{k}\to\operatorname{\mathcal{H}}$ that outputs a predictor $A(S)\in\operatorname{\mathcal{H}}$. Given a predictor $h$ and a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$, the empirical loss of the predictor over a set of $m$ independently and identically distributed (i.i.d.) examples is ^$\widehat{\mathcal{L}}_{S}(h)=\frac{1}{m}\sum_{j=1}^{m}\ell(h(\mathbf{x}_{j}),y_{j})$. We denote $\operatorname{\mathcal{D}}$ the data-generating distribution over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$ such that $S\sim\operatorname{\mathcal{D}}^{m}$ and the generalization loss of a predictor $h$ is $\mathcal{L}_{\mathcal{D}}(h)=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell(h(\mathbf{x}),y)\right]$.
@@ 2.1 PAC-Bayesian Learning Framework
For any distribution $\operatorname{\mathcal{D}}$ on $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any set $\operatorname{\mathcal{H}}$ of predictors $h:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$, for any loss $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to[0,1]$, for any prior distribution $P$ on $\operatorname{\mathcal{H}}$, for any $\delta\in(0,1]$ and for any convex function $\Delta:[0,1]\times[0,1]\to\operatorname{\mathbb{R}}$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, we have
@@
A common choice of comparator function is given by the Kullback-Leibler divergence between two Bernoulli distributions of probability of success $q$ and $p$,
@@ 2.2 Sample Compression Learning Framework
Initiated by Littlestone & Warmuth (1986) and refined by many authors over time (Attias et al., 2018; Bazinet et al., 2024; Campi & Garatti, 2023; David et al., 2016; Floyd & Warmuth, 1995; Hanneke & Kontorovich, 2021; Hanneke et al., 2018, 2019, 2024; Laviolette et al., 2005; Marchand & Shawe-Taylor, 2002; Marchand & Sokolova, 2005; Moran & Yehudayoff, 2016; Rubinstein & Rubinstein, 2012; Shah, 2007), the sample compression theory expresses generalization bounds on predictors that rely only on a small subset of the training set, referred to as the compression set, valid even if the learning algorithm observes the entire training dataset.The sample compression theorems thus express the generalization ability of predictive models as an accuracy-complexity trade-off, measured respectively by the training loss and the size of the compressed representation.
@@
Notable theoretical results. Theorem 2.2 below, due to Laviolette et al. (2005), improves the bound developed for the SCM algorithm (Marchand & Shawe-Taylor, 2001, 2002; Marchand & Sokolova, 2005). It is premised on two data-independent distributions: $P_{\mathscr{P}({\mathbf{m}})}$ on the compression set indices $\mathscr{P}({\mathbf{m}})$, and $P_{\Omega}$ on a discrete set of messages $\Omega$. Noteworthy, the bound is valid solely for binary-valued losses, as it considers each "successful" and "unsuccessful" prediction to be the result of a Bernoulli distribution.
@@
Theorem 2.2 is limited in its scope, for many tasks involve non-binary loss (e.g. regression tasks, cost-sensitive classification). The following recent result, due to Bazinet et al. (2024), permits real-valued losses $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to\operatorname{\mathbb{R}}$. Given a comparator function $\Delta$ analogous to the one classically used in PAC-Bayesian theorems, Theorem 2.3 bounds the discrepancy between the empirical loss of the reconstructed hypothesis $\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)$ on the complement set $S_{\mathbf{j}}$ and its generalization loss on the data distribution ${\mathcal{D}}$.
Theorem 2.3 (Sample compression - real-valued losses (Bazinet et al., 2024)).
@@
2.3 A New-PAC-Bayes Sample Compression Learning
Scheme
Our first contribution lies in the extension of Theorem 2.3 to real-valued messages. This is achieved by using a strategy from the PAC-Bayesian theory: we consider a data-independent prior distribution over the messages $\Omega$, denoted $P_{\Omega}$, and a data-dependent posterior distribution over the messages, denoted $Q_{\Omega}$. We then obtain a bound for the expected loss over $Q_{\Omega}$. Note that this new result shares similarities with the existing PAC-Bayes sample compression theory (Germain et al., 2011, 2015; Laviolette & Marchand, 2005, 2007), which gives PAC-Bayesian bounds for an expectation of data-dependent predictors given distributions on both the compression set and the messages. Our result differs by restricting the expectation solely according to the message.
@@ 2.3 A New-PAC-Bayes Sample Compression Learning
See Appendix D for the complete proof of Theorem 2.4 and specific choices of $\Delta$. In particular, ${\mathcal{A}}_{\mathrm{kl}}(m)\leq\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}2\sqrt{m-\operatorname{|\mathbf{j}|}}\leq 2\sqrt{m-\min_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}$ (see Corollary D.2).
Note that, in the setting where the compression set size is fixed (i.e., $\forall{\mathbf{j}}\in J,\ \operatorname{|\mathbf{j}|}=c$), Theorem 2.3 is a particular case of Theorem 2.4 where we restricted the message space $\Omega$ to be discrete and the posterior $Q_{\Omega}$ is a Dirac over a single $\omega^{*}\in\Omega$ (i.e., $Q_{\Omega}(\omega^{*})=1$).
Leveraging the PAC-Bayes disintegrated theorem of Viallard et al. (2024), we also obtain a variant of Theorem 2.4 valid for a single deterministic predictor $\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)$ (with $\omega$ drawn once according to $Q_{\Omega}$), instead of the expected loss according to $Q_{\Omega}$. This new result require the use of a singular compression function $\zeta(S,P_{\Omega})=(S_{\operatorname{\mathbf{j}}},Q_{\Omega}^{S})$, that takes as input a dataset and a prior distribution over the messages, and outputs a compression set and a posterior distribution.
@@
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any set $J\subseteq\operatorname{\mathscr{P}}(\mathbf{m})$, for any prior distribution $P_{J}$ over $J$, for any prior distribution $P_{\Omega}$ over $\Omega$, for any reconstruction function $\operatorname{\mathscr{R}}$, for any compression function $\zeta$, for any loss $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to[0,1]$, for any $\alpha>1$, for any convex function $\Delta:[0,1]\times[0,1]\to\operatorname{\mathbb{R}}$, for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, $\omega\sim Q_{\Omega}^{S}$, we have
$(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))$
$\leq\frac{2\alpha-1}{\alpha-1}\ln\frac{2}{\delta}+\ln\frac{1}{P_{J}(\operatorname{\mathbf{j}})}+D_{\alpha}(Q_{\Omega}^{S}||P_{\Omega})+\ln{\mathcal{A}}_{\Delta}(m)$
with $(S_{\operatorname{\mathbf{j}}},Q_{\Omega}^{S})=\zeta(S,P_{\Omega})$.
@@ 3 Meta-Learning with Hypernetworks
Let a task $\operatorname{\mathcal{D}}_{i}$ be a realization of a meta distribution ${\mathbf{D}}$, and $S_{i}{\sim}{\mathcal{D}}_{i}^{m_{i}}$ be a dataset of $m_{i}$ i.i.d. samples from a given task. A meta-learning algorithm receives as input a meta-dataset $\mathbf{S}=\{S_{i}\}_{i=1}^{n}$, that is a collection of $n$ datasets obtained from distributions $\{\operatorname{\mathcal{D}}_{i}\}_{i=1}^{n}$. The aim is to learn a meta-predictor $\operatorname{\mathscr{H}}$ that, given only a few sample $S^{\prime}\sim({\mathcal{D}}^{\prime})^{|S^{\prime}|}$ for a new task ${\mathcal{D}}^{\prime}\sim{\mathbf{D}}$, produces a predictor $h^{\prime}=\operatorname{\mathscr{H}}(S^{\prime})$ that generalizes well, i.e., with low generalization loss $\mathcal{L}_{\mathcal{D^{\prime}}}(h^{\prime})$.
@@
Toward generalization bounds. From now on, our aim is to study variants of hypernetwork architectures for the meta-predictor $\operatorname{\mathscr{H}}_{\theta}$, each variant inspired by the learning frameworks of Section 2. By doing so, once $\operatorname{\mathscr{H}}_{\theta}$ is learned, every downstream network $h_{\gamma^{\prime}}=\operatorname{\mathscr{H}}(S^{\prime})$ comes with its own risk certificate, i.e., an upper bound on its generalization loss ${\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})$ statistically valid with high probability.

### v3

@@ 1 Introduction
Machine learning is increasingly shaping our daily lives, a trend accelerated by rapid advancements in deep learning and the widespread deployment of these models across various applications (e.g., language models and AI agents). Ensuring the reliability of machine learning models is therefore more critical than ever. A fundamental aspect of reliability is understanding how well a model generalizes beyond its training data, particularly for modern neural networks, whose complexity makes it challenging to obtain strong theoretical guarantees. One way to assess generalization is through probabilistic bounds on a model’s error rate. Applying these techniques to deep neural networks is challenging because classical approaches struggle to account for the true effective complexity of such models (Zhang et al., 2017), which may not be well captured by naive measures such as parameter count (Belkin et al., 2019). However, prior work suggests that if one can obtain a more compact yet expressive representation of a model’s complexity, tighter generalization bounds are possible (Dziugaite et al., 2020; Wang et al., 2022; Kawaguchi et al., 2023), even when it comes to models as big as Large Langage Models (Lotfi et al., 2024a, b).
@@
Related works.  The meta-learning framework was pioneered by Baxter (2000), where a learning problem encompasses multiple tasks under the assumption that all learning tasks are independently and identically distributed (i.i.d.) from a task environment distribution. Various standpoints have been considered to derive generalization guarantees in this setting, for example the VC-theory (Maurer, 2009; Maurer et al., 2016) and algorithmic stability (Maurer, 2005; Chen et al., 2020). In recent years, the PAC-Bayesian framework has been the foundation of a rich line of work: Pentina & Lampert (2014); Amit & Meir (2018); Rothfuss et al. (2021); Liu et al. (2021); Guan et al. (2022); Guan & Lu (2022); Rezazadeh (2022); Zakerinia et al. (2024). Up to our knowledge, using the sample compression framework (Littlestone & Warmuth, 1986) to derive generalization bounds for meta-learned predictors is an idea that has not been explored yet.
@@ 2 Learning Theory and Generalization Bounds
The prediction problem.  A dataset $S=\{({\mathbf{x}}_{j},y_{j})\}_{j=1}^{m}$ is a collection of $m$ examples, each of them being a feature-target pair $(\mathbf{x},y)\in\mathcal{X}\times\mathcal{Y}$, and a predictor is a function $h:\mathcal{X}\rightarrow\mathcal{Y}$. We denote $\operatorname{\mathcal{H}}$ the predictor space. Let $A$ be a learning algorithm $A:\bigcup_{k\in\mathbb{N}}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{k}\to\operatorname{\mathcal{H}}$ that outputs a predictor $A(S)\in\operatorname{\mathcal{H}}$. Given a predictor $h$ and a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$, the empirical loss of the predictor over a set of $m$ i.i.d. examples is ^$\widehat{\mathcal{L}}_{S}(h)=\frac{1}{m}\sum_{j=1}^{m}\ell(h(\mathbf{x}_{j}),y_{j})$. We denote $\operatorname{\mathcal{D}}$ the data-generating distribution over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$ such that $S\sim\operatorname{\mathcal{D}}^{m}$ and the generalization loss of a predictor $h$ is $\mathcal{L}_{\mathcal{D}}(h)=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell(h(\mathbf{x}),y)\right]$.
@@ 2.1 PAC-Bayesian Learning Framework
For any distribution $\operatorname{\mathcal{D}}$ on $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any set $\operatorname{\mathcal{H}}$ of predictors $h:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$, for any loss $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to[0,1]$, for any prior distribution $P$ over $\operatorname{\mathcal{H}}$, for any $\delta\in(0,1]$ and for any convex function $\Delta:[0,1]\times[0,1]\to\operatorname{\mathbb{R}}$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, we have
@@
A common choice of comparator function is the Kullback-Leibler divergence between two Bernoulli distributions of probability of success $q$ and $p$,
@@ 2.2 Sample Compression Learning Framework
Initiated by Littlestone & Warmuth (1986) and refined by many authors over time (Attias et al., 2024; Bazinet et al., 2025; Campi & Garatti, 2023; David et al., 2016; Hanneke & Kontorovich, 2021; Hanneke et al., 2019, 2024; Laviolette et al., 2005; Marchand & Shawe-Taylor, 2002; Marchand & Sokolova, 2005; Moran & Yehudayoff, 2016; Rubinstein & Rubinstein, 2012; Shah, 2007), the sample compression theory expresses generalization bounds on predictors that rely only on a small subset of the training set, referred to as the compression set, valid even if the learning algorithm observes the entire training dataset. The sample compression theorems thus express the generalization ability of predictive models as an accuracy-complexity trade-off, measured respectively by the training loss and the size of the compressed representation.
@@
An example: the Support Vector Machine (SVM).  Consider $A$ to be the SVM building algorithm, and $S$ a dataset with $\mathcal{Y}=\{-1,+1\}$. An easy way of reconstructing $A(S)$ is by choosing the compression set to be $S_{\mathbf{j}}$, the $m$ support vectors of $A(S)$, and having the reconstruction function to be the linear SVM algorithm $R=A$. Thus, we know that $R(S_{\mathbf{j}})=A(S_{\mathbf{j}})=A(S)$, without having to use any message.
Another example: the example-aligned decision stump.
Given a dataset $S$ with $\mathcal{Y}=\{-1,+1\}$, the example-aligned decision stump (weak) learning algorithm $A$ returns a predictor $A(S)=h_{\mathbf{x}^{\prime},\upsilon,k}$ such that $h_{\mathbf{x}^{\prime},\upsilon,k}(\mathbf{x})=1$ if $\upsilon(x_{k}-x^{\prime}_{k})\leq 0$ or $-1$ otherwise, for some datapoint $\mathbf{x}^{\prime}=(x_{1}^{\prime},\dots,x_{d}^{\prime})\in S$, direction $\upsilon\in\{-1,+1\}$ and index $k\in\{1,\dots,d\}$. We can fully reconstruct the stump with $R(S_{\mathbf{j}},\upsilon)=h_{\mathbf{x}^{\prime},\upsilon,k}$, where $S_{\mathbf{j}}=\{(\mathbf{x}^{\prime},y^{\prime})\}$ is the compression set and $\omega=(\upsilon,k)$ is the message.
Notable theoretical results.  Theorem 2.2 below, due to Laviolette et al. (2005), improves the bound developed for their Set Covering Machine (SCM) algorithm (Marchand & Shawe-Taylor, 2001, 2002; Marchand & Sokolova, 2005). It is premised on two data-independent distributions: $P_{\mathscr{P}({\mathbf{m}})}$ on the compression set indices $\mathscr{P}({\mathbf{m}})$, and $P_{\Omega}$ on a discrete set of messages $\Omega$. Noteworthy, the bound is valid solely for binary-valued losses, as it considers each "successful" and "unsuccessful" prediction to be the result of a Bernoulli distribution.
@@
Theorem 2.2 is limited in its scope, for many tasks involve non-binary losses (e.g. regression tasks, cost-sensitive classification). The following recent result, due to Bazinet et al. (2025), permits real-valued losses $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to\operatorname{\mathbb{R}}$. Given a comparator function $\Delta$ analogous to the one classically used in PAC-Bayesian theorems, Theorem 2.3 bounds the discrepancy between the empirical loss of the reconstructed hypothesis $\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)$ on the complement set $S_{\overline{\operatorname{\mathbf{j}}}}$ and its generalization loss on the data distribution ${\mathcal{D}}$.
Theorem 2.3 (Sample compression - real-valued losses (Bazinet et al., 2025)).
@@
2.3 A New PAC-Bayes Sample Compression Framework
Our first contribution lies in the extension of Theorem 2.3 to real-valued messages. This is achieved by using a strategy from the PAC-Bayesian theory: we consider a data-independent prior distribution over the messages $\Omega$, denoted $P_{\Omega}$, and a data-dependent posterior distribution over the messages, denoted $Q_{\Omega}$. We then obtain a bound for the expected loss over $Q_{\Omega}$. Note that this new result shares similarities with the existing PAC-Bayes sample compression theory (Germain et al., 2011, 2015; Laviolette & Marchand, 2005), which gives PAC-Bayesian bounds for an expectation of data-dependent predictors given distributions on both the compression set and the messages. Our result differs by restricting the expectation solely according to the message.
@@ 2.3 A New PAC-Bayes Sample Compression Framework
See Appendix D for the complete proof of Theorem 2.4 and specific choices of $\Delta$. In particular, ${\mathcal{A}}_{\mathrm{kl}}(m){\leq}\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}2\sqrt{m{-}\operatorname{|\mathbf{j}|}}{\leq}2\sqrt{m{-}\!\min_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}$ (Cor. D.2). Note that, in the setting where the compression set size is fixed (i.e., $\forall{\mathbf{j}}\in J,\ \operatorname{|\mathbf{j}|}{=}c$), Theorem 2.3 is a particular case of Theorem 2.4 where the message space $\Omega$ is discrete and the posterior $Q_{\Omega}$ is a Dirac over a single $\omega^{\star}{\in}\Omega$ (i.e., $Q_{\Omega}(\omega^{\star}){=}1$).
Leveraging the PAC-Bayes disintegrated theorem of Viallard et al. (2024), we also obtain a variant of Theorem 2.4 valid for a single deterministic predictor $\operatorname{\mathscr{R}}(S_{\mathbf{j}^{\star}},\omega^{\star})$ (with $\omega^{\star}$ drawn once according to $Q_{\Omega}$), instead of the expected loss according to $Q_{\Omega}$. This new result requires the use of a singular compression function $\zeta(S,P_{\Omega})=(S_{\operatorname{\mathbf{j}}^{\star}},Q_{\Omega}^{S})$, which takes as input a dataset and a prior distribution over the messages, and outputs a compression set and a posterior distribution.
@@
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any set $J\subseteq\operatorname{\mathscr{P}}(\mathbf{m})$, for any prior distribution $P_{J}$ over $J$, for any prior distribution $P_{\Omega}$ over $\Omega$, for any reconstruction function $\operatorname{\mathscr{R}}$, for any compression function $\zeta$, for any loss $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to[0,1]$, for any $\alpha>1$, for any convex function $\Delta:[0,1]\times[0,1]\to\operatorname{\mathbb{R}}$, for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$ (which leads to $(S_{\operatorname{\mathbf{j}}^{\star}},Q_{\Omega}^{S})=\zeta(S,P_{\Omega})$ ), and $\omega^{\star}\sim Q_{\Omega}^{S}$, we have
$(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}^{\star}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})))$
$\leq\frac{2\alpha-1}{\alpha-1}\ln\!\frac{2}{\delta}+\ln\!\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}+D_{\alpha}(Q_{\Omega}^{S}||P_{\Omega})+\ln{\mathcal{A}}_{\Delta}(m)\,.$
@@ 3 Meta-Learning with Hypernetworks
Let a task $\operatorname{\mathcal{D}}_{i}$ be a realization of a meta distribution ${\mathbf{D}}$, and $S_{i}{\sim}{\mathcal{D}}_{i}^{m_{i}}$ be a dataset of $m_{i}$ i.i.d. samples from a given task. A meta-learning algorithm receives as input a meta-dataset $\mathbf{S}=\{S_{i}\}_{i=1}^{n}$, that is a collection of $n$ datasets obtained from distributions $\{\operatorname{\mathcal{D}}_{i}\}_{i=1}^{n}$. The aim is to learn a meta-predictor $\operatorname{\mathscr{H}}$ that, given only a few sample $S^{\prime}\sim({\mathcal{D}}^{\prime})^{m^{\prime}}$ for a new task ${\mathcal{D}}^{\prime}\sim{\mathbf{D}}$, produces a predictor $h^{\prime}=\operatorname{\mathscr{H}}(S^{\prime})$ that generalizes well, i.e., with low generalization loss $\mathcal{L}_{\mathcal{D^{\prime}}}(h^{\prime})$.
@@
Toward generalization bounds.  From now on, our aim is to study variants of hypernetwork architectures for the meta-predictor $\operatorname{\mathscr{H}}_{\theta}$, each variant inspired by the learning frameworks of Section 2. By doing so, once $\operatorname{\mathscr{H}}_{\theta}$ is learned, every downstream network $h_{\gamma^{\prime}}=\operatorname{\mathscr{H}}_{\theta}(S^{\prime})$ comes with its own risk certificate, i.e., an upper bound on its generalization loss ${\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})$ statistically valid with high probability.
We stress that, while the usual meta-learning bounds are computed after the meta-learning training phase in order to guarantee the expected performance of future downstream predictors learned on yet unseen tasks, the bounds we provide concern the generalization of downstream predictors once they are outputted by the meta-predictor.